OpenStudy (anonymous):

f(x)=2e^(-x),f'(X)=?

hartnn (hartnn):

\(\huge \color{blue}{\frac{d}{dx}e^{ax}=ae^{ax}}\)

hartnn (hartnn):

here a=-1 and 2can be factored out of derivative because its a constant.

hartnn (hartnn):

@saifoo.khan , you meant \(\Huge {2e^{-x}[}{\frac{d}{dx} (-x)}]\) right ?

OpenStudy (anonymous):

-2e is the ans

hartnn (hartnn):

now u know how to get \(-2e^{-x}\) as answer ?

OpenStudy (saifoo.khan):

Yes @hartnn . did a mistake. :D

OpenStudy (anonymous):

no i m not get it

hartnn (hartnn):

which part ?

OpenStudy (anonymous):

-x will remain as power also?

hartnn (hartnn):

yes, power will never change for e, it only gets multiplied before e term

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

f'x=(2e-x)'=(2/ex)'=(2'ex-2(ex)')/(ex)2 =(-2ex)/(ex)2 =-2e-x --i think it is the answer!!!

hartnn (hartnn):

this i got from chain rule, \(\huge \color{blue}{\frac{d}{dx}e^{ax}=ae^{ax}}\)

OpenStudy (anonymous):

Is the result u found is the same with me?

hartnn (hartnn):

yes.

OpenStudy (anonymous):

Thanks lol!